# Covering Multiple Objectives with a Small Set of Solutions Using Bayesian Optimization

### v1

@@ A.4 Additional Implementation Details
A.4.1 Trust Region Hyperparameters
For all trust region methods, the trust region hyperparameters are set to the TuRBO defaults used by Eriksson et al. (2019).
A.4.2 Surrogate Model
Since the tasks considered in this paper are challenging high-dimensional tasks, requiring a large number of function evaluations, we use approximate GP surrogate models. In particular, we use a PPGPR (Jankowiak et al., 2020) surrogate models with a constant mean, standard RBF kernel, and $1024$ inducing points. Additionally, we use a deep kernel (several fully connected layers between the search space and the GP kernel) (Wilson et al., 2016). We use two fully connected layers with $D$ nodes each, where $D$ is the dimensionality of the search space.
We use the same PPGPR model(s) with the same configuration for MOCOBO, TuRBO, LOL-BO, and ROBOT. For MOCOBO, to model the $T$-dimensional output space, we use $T$ PPGPR models, one to approximate each objective $f_{1},\ldots,f_{T}$. To allow information sharing between the models, we use a shared deep kernel (the $T$ PPGPR models share share the same two-layer deep kernel) (Jiang et al., 2018; Patacchiola et al., 2019).
Unlike the other methods compared, MORBO was designed for use with an exact GP model rather than an approximate GP surrogate model. For fair comparison, we therefore run MORBO with an exact GP using all default hyperparameters and the official codebase provided by Daulton et al. (2021).
We train the PPGPR surrogate model(s) on data collected during optimization using the Adam optimizer with a learning rate of $0.001$ and a mini-batch size of $256$. On each step of optimization, we update the model on collected data until we stop making progress (loss stops decreasing for $3$ consecutive epochs), or exceed $30$ epochs. Since we collect a large amount of data for each optimization run (i.e. as many as $62e$ 6 data points in a single for the “template constrained” peptide design task), we avoid updating the model on all data collected at once. On each step of optimization, we update the current surrogate model only on a subset of $1000$ of the collected data points. This subset is constructed of the data that has obtained the highest objective values so far, along with the most recent batch of data collected. By always updating on the most recent batch of data collected, we ensure that the surrogate model is conditioned on every data point collected at some point during the optimization run.
A.4.3 Initialization Data
@@
To initialize optimization for the molecule design task, we take a random subset of $10000$ molecules from the standardized unlabeled dataset of 1.27M molecules from the Guacamol benchmark software (Brown et al., 2019). We generate labels for these $10,000$ molecules once, and then use the labeled data to initialize optimization for all methods compared.
@@
A.4.4 Diversity Constraints and Associated Hyperparameters for ROBOT Baseline
For a single objective, ROBOT seeks a diverse set of $M$ solutions, requiring that the set of solutions have a minimum pairwise diversity $\tau$ according to the user specified diversity function $\delta{}$. Since Maus et al. (2023) also consider rover and molecule design tasks, we use the same diversity function $\delta{}$ and diversity threshold $\tau$ used by Maus et al. (2023) for these two tasks. For the peptide design tasks, we define $\delta{}$ to be the edit distance between peptide sequences, and use a diversity threshold of $\tau=3$ edits. For the image optimization task, since there is no obvious semantically meaningful measure of diversity between two sets of input parameters, we define $\delta{}$ to be the euclidean distance between solutions, and use $\tau=1.45$, the approximate average euclidean distance between a randomly selected pair of points in the search space.
A.5 Additional Task Details
@@
A.5.1 Additional Details for Image Tone Mapping Task
@@
1 nima Talebi & Milanfar,2018
2 nima-vgg16-ava Talebi & Milanfar,2018; Murray et al.,2012
3 topiq-iaa-res50 Chen et al.,2024
4 laion-aes Schuhmann et al.,2022
5 hyperiqa Su et al.,2020
6 tres Golestaneh et al.,2022
7 liqe Zhang et al.,2023
Table A.3: pyiqa metric ID strings used to identify the $T=7$ target image quality metrics used for the image tone mapping task. Each metrics is a no-reference image aesthetic (IAA) and quality (IQA) assessment metric obtained from the pyiqa library (Chen & Mo, 2022).
@@
We use two benchmark images. The first is the “Stanford Memorial Church” image obtained from https://www.pauldebevec.com/Research/HDR/ by courtesy of Paul E. Debevec (Debevec & Malik, 1997). The second is the “desk lamp” image obtained from https://cadik.posvete.cz/tmo/ by courtesy of Martin Čadík (Čadík, 2008).
@@
We consider a tone mapping pipeline consisting of a multi-layer detail decomposition (Tumblin & Turk, 1999; Durand & Dorsey, 2002; Li et al., 2005) using the guided filter by He et al. (2013) (3 detail layers and 1 base layer), followed by gamma correction (Reinhard et al., 2005, Section 2.9). The resulting optimization problem is 13-dimensional.
The tone mapping pipeline is very similar to the classic approach proposed by Tumblin & Turk (1999). the main difference is that, similarly to Farbman et al. (2008), we replace the diffusion smoothing filter with a more recent edge-preserving detail smoothing filter, the guided filter, by He et al. (2013). (See a similar approach by Farbman et al. (2008).) First, given an HDR image in the RGB color space $I=(I_{\mathrm{r}},I_{\mathrm{r}},I_{\mathrm{b}}))$, where $I_{\mathrm{r}}$ is the red channel, $I_{\mathrm{r}}$ is the blue channel, and $I_{\mathrm{g}}$ is the green channel, we compute the luminance according to
@@
(The constants were taken from the code of Farbman et al. 2008.) The luminance image $L$ is then logarithmically compressed and then decomposed into three detail layers and one base layer by applying the guided filter three times, each with a different radius parameter $r_{i}$ and smoothing parameter $\epsilon_{i}$ for $i=1,\ldots,3$. Then, we amplify or attenuate each channel with a corresponding gain coefficient $g_{\mathrm{detail},i}$ $i=1,\ldots,3$ for the detail and $g_{\mathrm{base}}$ for the base layers. The image is then reconstructed by adding all the layers, including the colors. Following Tumblin & Turk (1999), we also apply a gain, $g_{\mathrm{color}}$, to the color channels. Finally, the resulting image is applied an overall gain $g_{\mathrm{out}}$ and then gamma-corrected (Reinhard et al., 2005, Section 2.9) with an exponent of $1/\gamma$. The parameters for this pipeline are organized in Table A.4. The implementation uses OpenCV (Bradski, 2000), in particular, the guided filter implementation in the extended image processing (ximgproc) submodule.
@@
A.5.2 Additional Details for Peptide Design Task
@@
Table A.6: Template amino acid sequences used for the “template constrained” peptide design task.
Table A.5 specifies the $T=11$ target bacteria used for the peptide design task from Section 4. The first seven bacteria are Gram negative bacteria (Objective IDs B1-B7) and the last four (Objective IDs B8-B11) are Gram positive. Table A.6 gives the $10$ template amino acid sequences used for the “template constrained” variation of the peptide design task from Section 4.
A.6 Proof that Finding $S^{*}_{D_{s}}$ is NP-Hard
@@
  • A collection of $n$ subsets $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{n}\}$, where $S_{i}\subseteq U$.
@@
The objective is to find a collection of $K$ subsets $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ such that the total number of elements covered, $\bigcup_{S\in\mathcal{S}^{\prime}}S$, is maximized.
@@
    Let each subset $S_{i}\in\mathcal{S}$ correspond to a point $\mathbf{x}_{i}\in D_{s}$.
@@
    Define $f_{t}(\mathbf{x}_{i})=1$ if subset $S_{i}$ covers element $e_{t}$, and $f_{t}(\mathbf{x}_{i})=0$ otherwise.
@@
  • Each subset $S_{i}\in\mathcal{S}$ corresponds to a point $\mathbf{x}_{i}\in D_{s}$.
  • If subset $S_{i}$ covers element $e_{t}$, then $f_{t}(\mathbf{x}_{i})=1$; otherwise, $f_{t}(\mathbf{x}_{i})=0$.
@@
A.7 Proofs for Greedy Approximation Algorithm for Finding $S^{*}_{D_{s}}$
In this section, we prove Theorem 3.2, and Corollary 3.3. Proofs will make use of the following definitions:
@@ A.7 Proofs for Greedy Approximation Algorithm for Finding $S^{*}_{D_{s}}$
A.7.1 Approximation Proof for Algorithm 1
In this section, we prove Theorem 3.2.
Proof That Algorithm 1 Is a $(1-\frac{1}{e})$-Approximation
@@ A.7.1 Approximation Proof for Algorithm 1
$\max_{\mathbf{x}\in S^{\prime}}f_{t}(\mathbf{x})=\max\left(\max_{\mathbf{x}\in S}f(x),\max_{\mathbf{x}\in S^{\prime}\setminus S}\right)\geq\max_{\mathbf{x}\in S}f(x).$
@@
$c(S\cup\{x^{*}\})-c(S)$ $=\sum_{i=1}^{T}\left(\max_{x\in S\cup\{x^{*}\}}f_{i}(x)-\max_{x\in S}f_{i}(x)\right)$
$=\sum_{i=1}^{T}\max\left[f_{i}(x^{*}),\max_{\mathbf{x}\in S}f_{i}(x)\right]-\max_{\mathbf{x}\in S}f_{i}(x)$
$=\sum_{i=1}^{T}\begin{cases}f_{i}(\mathbf{x}^{*})-\max_{\mathbf{x}\in S}f_{i}(\mathbf{x})&f_{i}(\mathbf{x}^{*})>\max_{\mathbf{x}\in S}f_{i}(\mathbf{x})\\ 0&\textrm{otherwise}\end{cases}$
@@
$c(S^{\prime}\cup\{x^{*}\})-c(S^{\prime})=\sum_{i=1}^{T}\begin{cases}f_{i}(\mathbf{x}^{*})-\max_{\mathbf{x}\in S^{\prime}}f_{i}(\mathbf{x})&f_{i}(\mathbf{x}^{*})>\max_{\mathbf{x}\in S^{\prime}}f_{i}(\mathbf{x})\\ 0&\textrm{otherwise}\end{cases}$
@@
$\max_{\mathbf{x}\in S^{\prime}}f_{i}(\mathbf{x})=\max\left[\max_{\mathbf{x}\in S}f_{i}(\mathbf{x}),\max_{\mathbf{x}\in S^{\prime}\setminus S}f_{i}(\mathbf{x})\right]\geq\max_{\mathbf{x}\in S}f_{i}(\mathbf{x}),$
we know then that $f_{i}(\mathbf{x}^{*})-\max_{\mathbf{x}\in S^{\prime}}\leq f_{i}(\mathbf{x}^{*})-\max_{\mathbf{x}\in S}f_{i}(\mathbf{x})$, and further that $f_{i}(\mathbf{x}^{*})>\max_{\mathbf{x}\in S^{\prime}}f_{i}(\mathbf{x})$ implies that $f_{i}(\mathbf{x}^{*})>\max_{\mathbf{x}\in S}f_{i}(\mathbf{x})$. Taken together, these imply that:
$\sum_{i=1}^{T}\begin{cases}f_{i}(\mathbf{x}^{*})-\max_{\mathbf{x}\in S^{\prime}}f_{i}(\mathbf{x})&f_{i}(\mathbf{x}^{*})>\max_{\mathbf{x}\in S^{\prime}}f_{i}(\mathbf{x})\\ 0&\textrm{otherwise}\end{cases}\leq\sum_{i=1}^{T}\begin{cases}f_{i}(\mathbf{x}^{*})-\max_{\mathbf{x}\in S}f_{i}(\mathbf{x})&f_{i}(\mathbf{x}^{*})>\max_{\mathbf{x}\in S}f_{i}(\mathbf{x})\\ 0&\textrm{otherwise}\end{cases}$
@@
This is a well know result about monotone submodular functions, show for example by Nemhauser et al. (1978). ∎

### v2

@@ A.7 Additional implementation details
A.7.1 Trust region hyperparameters
For all trust region methods, the trust region hyperparameters are set to the TuRBO defaults used by Eriksson et al. [17].
A.7.2 Surrogate model
Since the tasks considered in this paper are challenging high-dimensional tasks, requiring a large number of function evaluations, we use approximate GP surrogate models. In particular, we use a PPGPR [28] surrogate models with a constant mean, standard RBF kernel, and $1024$ inducing points. Additionally, we use a deep kernel (several fully connected layers between the search space and the GP kernel) [70]. We use two fully connected layers with $D$ nodes each, where $D$ is the dimensionality of the search space.
We use the same PPGPR model(s) with the same configuration for MOCOBO, TuRBO, LOL-BO, and ROBOT. For MOCOBO, to model the $T$-dimensional output space, we use $T$ PPGPR models, one to approximate each objective $f_{1},\ldots,f_{T}$. To allow information sharing between the models, we use a shared deep kernel (the $T$ PPGPR models share share the same two-layer deep kernel) [29, 49].
Unlike the other methods compared, MORBO was designed for use with an exact GP model rather than an approximate GP surrogate model. For fair comparison, we therefore run MORBO with an exact GP using all default hyperparameters and the official codebase provided by Daulton et al. [11].
We train the PPGPR surrogate model(s) on data collected during optimization using the Adam optimizer with a learning rate of $0.001$ and a mini-batch size of $256$. On each step of optimization, we update the model on collected data until we stop making progress (loss stops decreasing for $3$ consecutive epochs), or exceed $30$ epochs. Since we collect a large amount of data for each optimization run (i.e. as many as $62e$ 6 data points in a single for the “template constrained" peptide design task), we avoid updating the model on all data collected at once. On each step of optimization, we update the current surrogate model only on a subset of $1000$ of the collected data points. This subset is constructed of the data that has obtained the highest objective values so far, along with the most recent batch of data collected. By always updating on the most recent batch of data collected, we ensure that the surrogate model is conditioned on every data point collected at some point during the optimization run.
A.7.3 Initialization data
@@
To initialize optimization for the molecule design task, we take a random subset of $10000$ molecules from the standardized unlabeled dataset of 1.27M molecules from the Guacamol benchmark software [6]. We generate labels for these $10,000$ molecules once, and then use the labeled data to initialize optimization for all methods compared.
@@
A.7.4 Diversity constraints and associated hyperparameters for ROBOT baseline
For a single objective, ROBOT seeks a diverse set of $M$ solutions, requiring that the set of solutions have a minimum pairwise diversity $\tau$ according to the user specified diversity function $\delta{}$. Since Maus et al. [45] also consider rover and molecule design tasks, we use the same diversity function $\delta{}$ and diversity threshold $\tau$ used by Maus et al. [45] for these two tasks. For the peptide design tasks, we define $\delta{}$ to be the edit distance between peptide sequences, and use a diversity threshold of $\tau=3$ edits. For the image optimization task, since there is no obvious semantically meaningful measure of diversity between two sets of input parameters, we define $\delta{}$ to be the euclidean distance between solutions, and use $\tau=1.45$, the approximate average euclidean distance between a randomly selected pair of points in the search space.
A.8 Additional task details
@@
A.8.1 HDR image tone mapping
@@
1 nima 59
2 nima-vgg16-ava 59, 46
3 topiq-iaa-res50 10
4 laion-aes 53
5 hyperiqa 58
6 tres 22
7 liqe 72
Table A.5: pyiqa metric ID strings used to identify the $T=7$ target image quality metrics used for image tone mapping. Each metric is a no-reference image aesthetic (IAA) or quality (IQA) assessment metric obtained from the pyiqa library [9].
@@
We use two benchmark images. The first is the “Stanford Memorial Church" image obtained from https://www.pauldebevec.com/Research/HDR/ by courtesy of Paul E. Debevec [12]. The second is the “desk lamp" image obtained from https://cadik.posvete.cz/tmo/ by courtesy of Martin Čadík [7]. Because commonly used metrics are only correlated with subjective image quality, prior work on tuning parameters for these and related benchmarks has been done by trial and error and human-in-the-loop type schemes [41], including preferential BO-based approaches [35, 36].
@@
We consider a tone mapping pipeline consisting of a multi-layer detail decomposition [62, 14, 38] using the guided filter by He et al. [25] (3 detail layers and 1 base layer), followed by gamma correction [51, Section 2.9], resulting in a 13-dimensional optimization problem. The complete image processing pipeline is very similar to the classic approach proposed by Tumblin and Turk [62], where the main difference is that, similarly to Farbman et al. [18], we replace the diffusion smoothing filter with a more recent edge-preserving detail smoothing filter, the guided filter, by He et al. [25]. First, given an HDR image in the RGB color space $I=(I_{\mathrm{r}},I_{\mathrm{r}},I_{\mathrm{b}}))$, where $I_{\mathrm{r}}$ is the red channel, $I_{\mathrm{r}}$ is the blue channel, and $I_{\mathrm{g}}$ is the green channel, we compute the luminance according to
@@
(The constants were taken from the code of Farbman et al. 18.) The luminance image $L$ is then logarithmically compressed and then decomposed into three detail layers and one base layer by applying the guided filter three times, each with a different radius parameter $r_{i}$ and smoothing parameter $\epsilon_{i}$ for $i=1,\ldots,3$. Then, we amplify or attenuate each channel with a corresponding gain coefficient $g_{\mathrm{detail},i}$ $i=1,\ldots,3$ for the detail and $g_{\mathrm{base}}$ for the base layers. The image is then reconstructed by adding all the layers, including the colors. Following Tumblin and Turk [62], we also apply a gain, $g_{\mathrm{color}}$, to the color channels. Finally, the resulting image is applied an overall gain $g_{\mathrm{out}}$ and then gamma-corrected [51, Section 2.9] with an exponent of $1/\gamma$. The parameters for this pipeline are organized in Table A.6. The implementation uses OpenCV [5], in particular, the guided filter implementation in the extended image processing (ximgproc) submodule.
@@
A.8.2 Peptide design
@@
Table A.8: Template amino acid sequences used for the “template constrained" peptide design task.
Table A.7 specifies the $T=11$ target bacteria used for the peptide design task from Section 4. The first seven bacteria are Gram negative bacteria (Objective IDs B1-B7) and the last four (Objective IDs B8-B11) are Gram positive. Table A.8 gives the $10$ template amino acid sequences used for the “template constrained" variation of the peptide design task from Section 4.
A.9 Proof that finding $S^{*}_{D_{s}}$ is NP-hard
@@
  • A collection of $n$ subsets $\mathcal{S}=\{A_{1},A_{2},\ldots,A_{n}\}$, where $A_{i}\subseteq U$.
@@
The objective is to find a collection of $K$ subsets $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ such that the total number of elements covered, $\bigcup_{A\in\mathcal{S}^{\prime}}A$, is maximized.
@@
    Let each subset $A_{i}\in\mathcal{S}$ correspond to a point $\mathbf{x}_{i}\in D_{s}$.
@@
    Define $f_{t}(\mathbf{x}_{i})=1$ if subset $A_{i}$ covers element $e_{t}$, and $f_{t}(\mathbf{x}_{i})=0$ otherwise.
@@
  • Each subset $A_{i}\in\mathcal{S}$ corresponds to a point $\mathbf{x}_{i}\in D_{s}$.
  • If subset $A_{i}$ covers element $e_{t}$, then $f_{t}(\mathbf{x}_{i})=1$; otherwise, $f_{t}(\mathbf{x}_{i})=0$.
@@
Proof significance
We do not claim the above proof of Lemma 3.1 as a novel contribution of this work, as it follows straightforwardly from the fact that the well-known Maximum Coverage Problem is NP-hard. Note that proving Lemma 3.1 would also follow straightforwardly from the work of Ding et al. [13] who demonstrate that the coverage optimization problem generalizes k-means clustering. Rather than providing additional novel contribution, the above proof of Lemma 3.1 serves to justify our use of a greedy approximation algorithm (Algorithm 1) to approximate $S^{*}_{D_{s}}$ on each iteration of MOCOBO.
A.10 Approximation proof for Algorithm 1
In this section, we prove Theorem 3.2, that Algorithm 1 is a $(1-\frac{1}{e})$-Approximation.
@@ A.10 Approximation proof for Algorithm 1
Definition A.5 (Contribution of Objective Function $f_{t}$ to Coverage Score).
Given a covering set $S\subseteq D_{s}$, we denote the contribution of objective $f_{t}\in\{f_{1},f_{2},\ldots,f_{T}\}$ to the overall coverage score $c(S)$ as $g_{t}(S)$ where
$g_{t}(S)=\max_{\mathbf{x}\in S}f_{t}(\mathbf{x})$
It follows that from Definition A.5 that:
$c(S)=\sum_{t=1}^{T}g_{t}(S)$
@@
$g_{t}(S^{\prime})=\max\left(g_{t}(S),g_{t}(S^{\prime}\setminus S)\right)\geq g_{t}(S).$
@@
$c(S\cup\{x^{*}\})-c(S)$ $=\sum_{t=1}^{T}\left(g_{t}(S\cup\{x^{*}\})-g_{t}(S)\right)$
$=\sum_{t=1}^{T}\max\left[f_{t}(x^{*}),g_{t}(S)\right]-g_{t}(S)$
$=\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S)&f_{t}(\mathbf{x}^{*})>g_{t}(S)\\ 0&\textrm{otherwise}\end{cases}$
@@
$c(S^{\prime}\cup\{x^{*}\})-c(S^{\prime})=\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S^{\prime})&f_{t}(\mathbf{x}^{*})>g_{t}(S^{\prime})\\ 0&\textrm{otherwise}\end{cases}$
@@
$g_{t}(S^{\prime})=\max\left[g_{t}(S),g_{t}(S^{\prime}\setminus S)\right]\geq g_{t}(S),$
we know then that $f_{t}(\mathbf{x}^{*})-g_{t}(S^{\prime})\leq f_{t}(\mathbf{x}^{*})-g_{t}(S)$, and further that $f_{t}(\mathbf{x}^{*})>g_{t}(S^{\prime})$ implies that $f_{t}(\mathbf{x}^{*})>g_{t}(S)$. Taken together, these imply that:
$\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S^{\prime})&f_{t}(\mathbf{x}^{*})>g_{t}(S^{\prime})\\ 0&\textrm{otherwise}\end{cases}\leq\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S)&f_{t}(\mathbf{x}^{*})>g_{t}(S)\\ 0&\textrm{otherwise}\end{cases}$
@@
This is a well know result about monotone submodular functions, show for example by Nemhauser et al. [48]. ∎